# On Nanocones as a Gravitational Analog System

### v1

@@ Abstract
This study delves into the fundamental properties of graphene and boron nitride (BN) nanostructures, exploring their torsional energy characteristics within the framework of Teleparallel Equivalent of General Relativity (TEGR). By constructing nanocones with disclination defects in these materials, we investigate the linear dependence of torsional energy on the disclination angle, as predicted by TEGR. The qualitative validation of TEGR’s energy expression is supported by our simulations, which show a strong correlation between the torsional energy and the disclination angle, consistent with the theoretical predictions. Furthermore, we propose a quantitative analysis by estimating the coupling constant $k$ associated with TEGR through molecular simulations and Density Functional Theory (DFT) calculations. Our results suggest that $k$ reflects the interatomic forces within the materials, providing insights into the nature of spacetime and gravitational interactions on a microscopic scale. These findings not only contribute to our understanding of material physics but also offer implications for the precision and validity of TEGR in describing gravitational phenomena.
June 8, 2024
@@ I Introduction
Analog models in physics have long existed. A mathematical equivalence between the equations governing certain physical phenomena allows for analogy between them. A good example of this occurred in the realm of electrical circuits. In the early 19th century, when instrumentation for measuring electrical and magnetic quantities was being developed, it was very costly to conduct investigations in circuits containing capacitors and inductors. Therefore, mechanical analog models were employed, which facilitated a better understanding of electrical circuits, as well as simpler experimentation. Analog models thus enable the testing of physical concepts without the need to access the physical system of interest. Experimental black hole physics is still in its infancy and could greatly benefit from the use of gravitational analog models barcelo . While experimenting with gravitational waves can shed light on certain aspects of black hole physics, testing other concepts such as Hawking radiation requires the use of analog models, often established in condensed matter. However, there are also acoustic analogs; acoustic black holes, proposed by Unruh Unruh , have been used to understand how event horizons can produce a spectrum of thermal radiation Unruh1 . Hence, gravitational analog models have become important tools for understanding black hole physics. One such example was the construction of an experimental apparatus where a robophysical model was built to simulate the motion of particles around a Schwarzschild singularity li2023robophysical , an experiment that culminated in very satisfactory results compared to the expected values from GR.
@@
General Gelativity (GR) is the most widely accepted theory in the literature describing gravitation. However, it is fraught with inconsistencies. Perhaps the most important of these is the definition of gravitational energy. Indeed, there are several attempts to define such an important concept in physics within the scope of general relativity, but none of them meets the necessary requirements for a consistent definition, namely: i) the energy should be independent of the coordinate system used; ii) it should be the component of a covariant four-vector under global Lorentz transformations and therefore should depend on the choice of reference frame. Alternatively, there are approaches that describe the gravitational field, in which the concept of energy is more natural. In this article, an alternative theory dynamically equivalent to general relativity, called Teleparallel Equivalent to General Relativity (TEGR) maluf , will be used. In the context of TEGR, there exists a consistent energy-momentum four-vector, so this quantity can be used to experimentally determine which approach correctly describes gravitation. In other words, from a dynamic point of view, both GR and TEGR predict the same behaviors. On the other hand, a consistent definition of energy exists only in TEGR (for a complete discussion of gravitational energy, we encourage the reader to consult Ref. annalen ). On one hand, there are currently no experiments designed to deal with gravitational energy. On the other hand, the use of gravitational analogue models may lead to clues suggesting the supremacy of TEGR over GR, if gravitational energy is confirmed in such systems.
@@
Graphene is a material predicted in the mid-20th century and synthesized in the early 21st century grafeno . It consists of a two-dimensional lattice of carbon atoms arranged in a hexagonal shape, giving graphene highly desired properties such as good conductivity and strength. The equation governing the dynamics of graphene is relativistic, as the Dirac equation embodies Lorentz transformations. The Dirac equation also naturally couples to the dynamic variables of TEGR (Teleparallel Equivalent of General Relativity), making graphene an excellent analog gravitational system that could allow investigation into the physical reality of gravitational energy.
@@
In pursuit of our goal, we consider a graphene sheet as a background flat geometry and impose a gravitational field by inserting a topological defect into the sheet. We choose a disclination defect to construct a well-known nanostructure for graphene, i.e., a nanocone charlier2001electronic ; adhikari2012stabilities ; ardeshana2017approach . Consequently, we evaluate the gravitational energy of the disclination using TEGR and compare the predicted qualitative behavior of the torsional energy with the energy associated with the torsion potential obtained from computational simulations within Universal Force Field (UFF) . The Universal Force Field (UFF) is a versatile molecular mechanics force field used in computational chemistry. It employs empirical parameters to model interatomic interactions, including bonded and non-bonded terms, torsional potentials, and geometric optimizations. UFF is widely applied in molecular simulations for accurate predictions of molecular structures, energies, and properties. The main aspects of this theory necessary for the comprehension of the simulations will be presented in this article. We also infer about the possibility of the physical origin of the coupling constant in graphene and obtain a numerical value for the energy within the order of magnitude given by the simulations. Therefore, we aim to address these aspects in this article to support future experiments designed to measure gravitational energy, especially since the synthesis of nanocones is already a reality experimental .
@@ II Teleparallelism Equivalent to General Relativity (TEGR)
Teleparallelism Equivalent to General Relativity (TEGR) is a geometrical theory of gravity that was first considered by Einstein in an attempt to unify gravity and electromagnetism. Several decades after being abandoned by Einstein, Møller revived the idea with another objective: to establish a true energy-momentum for the gravitational field moller1964conservation . Despite the many decades of GR, researchers have not been able to define a true energy-momentum tensor for the gravitational field, with several pseudo-tensors defined in the literature, all of which depend on the choice of coordinates. Therefore, a definition for the energy of a gravitational field is only possible globally, i.e., by integrating the pseudo-tensors across the entire spacetime. Such integration is possible only if the spacetime is asymptotically flat at spatial infinity. After a few years of attempting to define a true energy-momentum tensor for the gravitational field, Møller realized that the main problem with GR was the fact that its fundamental variable is the metric tensor $g_{\mu\nu}$ and concluded that he could solve the problem by formulating GR with the tetrad field $e_{a\mu}$ as its fundamental variable. Furthermore, Maluf realized that the field equations of the theory allow for a definition of a true energy-momentum tensor for the gravitational field. In the following, we follow the development of TEGR as presented in Ref. maluf .
@@
The geometry in which TEGR is established in the Weitzenböck geometry. Given the prominence in the literature of topological defects considered within Riemann-Cartan geometry, we depart from Riemann-Cartan to the Weitzenböck one, where the Levi-Civita connection is responsible for the rotation symmetry instead of the affine connection in Riemann-Cartan geometry. Therefore, we shall show that TEGR admits linear rotations (disclinations) and translations (dislocations) as topological defects in its geometry, similar to Riemann-Cartan.
In a manifold endowed with metricity, the structure equations can be obtained from the co-frame 1-form $e^{a}=e^{a}\,_{\mu}dx^{\mu}$ and the connection 1-form $\omega^{a}\,_{b}=\omega^{a}\,_{\mu b}dx^{\mu}$, where $\omega^{\lambda}\,_{\mu\nu}=e_{a}\,^{\lambda}e^{b}\,_{\nu}\omega^{a}\,_{\mu b}$ are the 0-form components of the connection. The torsion 2-form $T^{a}$ of the manifold can be obtained by the exterior total derivative $D$ of the co-vector, i.e.,
@@
$=\frac{1}{2}e^{a}\,_{\lambda}T^{\lambda}\,_{\mu\nu}dx^{\mu}\wedge dx^{\nu}$ (2)
@@
$T^{\lambda}\,_{\mu\nu}=e^{a}\,_{\lambda}(\partial_{\mu}e^{a}\,_{\nu}-\partial_{\nu}e^{a}\,_{\mu}+\omega^{\lambda}\,_{\mu\nu}-\omega^{\lambda}\,_{\nu\mu})$ (3)
@@
$=\frac{1}{2}e^{a}\,_{\alpha}e_{b}\,^{\beta}R^{\alpha}\,_{\beta\mu\nu}dx^{\mu}\wedge dx^{\nu}\,,$ (4)
@@
$R^{\alpha}\,_{\beta\mu\nu}$ $=\partial_{\mu}\omega^{\alpha}\,_{\nu\beta}-\partial_{\nu}\omega^{\alpha}\,_{\mu\beta}$
$+\omega^{\alpha}\,_{\mu\rho}\omega^{\rho}\,_{\nu\beta}-\omega^{\alpha}\,_{\nu\rho}\omega^{\rho}\,_{\mu\beta}$ (5)
@@
$\nabla_{\mu}e_{a\nu}=\partial_{\mu}e_{a\nu}-\Gamma^{\lambda}\,_{\mu\nu}e_{\lambda a}=0\,.$ (6)
@@
$\Gamma_{\mu\lambda\nu}=e^{a}\,_{\mu}\partial_{\lambda}e_{a\nu}$ (7)
@@
$T^{a}\,_{\lambda\nu}=\partial_{\lambda}e^{a}\,_{\nu}-\partial_{\nu}e^{a}\,_{\lambda}\,.$ (8)
@@
$K_{\mu\lambda\nu}$ $=$ $\frac{1}{2}(T_{\lambda\mu\nu}+T_{\nu\lambda\mu}+T_{\mu\lambda\nu})\,,$ (9)
can be evaluated. Since the affine connection is chosen to be zero, the components of the Levi-Civita connection $\mathring{\omega}_{\mu ab}$ can be evaluated using the relation
$\Gamma_{\mu\lambda\nu}=\mathring{\omega}_{\mu\lambda\nu}+K_{\mu\lambda\nu}=0$ (10)
@@
$\mathfrak{L}(e_{a\mu})$ $=$ $-\kappa\,e\Sigma^{abc}T_{abc}-\mathfrak{L}_{M}$ (12)
@@
-- $\mathfrak{L}_{M}\,,$
where $e=det(e^{a}\,_{\mu})$, $k=1/16\pi$ is the coupling constant ($k=c^{3}/16\pi G$ in SI units), and $\mathfrak{L}_{M}$ stands for the matter-radiation fields Lagrangian density. The torsion invariant, and hence the Lagrangian density, is equal to the curvature scalar constructed from the Levi-Civita connection plus a total divergence term, i.e.,
@@
Although the Lagrangian density (12) is not invariant under local Lorentz transformations, the field equations (14) are covariant under coordinate transformations and under the full group of Lorentz transformations. It can be shown that (14) may be written as $R_{a\mu}-\frac{1}{2}e_{a\mu}R=\frac{1}{2k}T_{a\mu}$, i.e., they are equivalent to Einstein’s equations. Therefore, TEGR yields the same dynamics as GR but allows for the definition of a true energy-momentum tensor.
@@ III Topological Defects in spacetime
The matrix $\mathcal{G}^{(i)(j)}$, called Frank matrix, represents the rotation matrix along the axis containing the angle deficit, i.e., the rotation angle represents the amount that we must rotate to perform a full $2\pi$ lap in spacetime. The curvature tensor of the affine connection is zero for the Weitzenböck geometry. However, the curvature tensor of the Levi-Civita connection $\mathring{\omega}_{\mu ab}$ is nonzero in the general case. This curvature tensor allows for a consistent definition of disclination in TEGR. This feature will be demonstrated here for the defect considered.
A conical defect along the $z$ direction is characterized by the deficit angle parameter $\Delta\phi=2\pi\big{(}1-\beta(r)\big{)}$ in the line element
@@
When $\beta=1$, we recover Minkowski spacetime. The deficit angle can be evaluated from (24) by computing the length of the circumference along the $z$ axis, but we will consider relations (22) and (23). First, we need a set of tetrads associated with the metric (24) and adapted to a static observer. One set satisfying the condition is given by
@@
$\mathring{\omega}_{2}\,^{(1)(2)}=1-\partial_{r}(r\,\beta)\,.$ (27)
@@
i.e., the rotation matrix of the deficit angle $\Delta\phi=2\pi(1-\beta_{0})$ along the $z$ axis. Thus, TEGR allows torsion and curvature and is consistent in computing topological defects in spacetime, with the rotational defects given by the Levi-Civita connection constructed out of the contortion tensor given by (10).
@@
Hence, the energy contained within a cylinder of radius $R>a$ and length $L$ is given by
$P^{(0)}$ $=2k\int_{V}\partial_{r}\big{(}1-\partial_{r}(r\beta)\big{)}dV$
@@
$=4\pi\,k\,L(1-\beta_{0})$
$=2k\,L\,\Delta\phi\,.$ (39)
From (39), we see that with TEGR, we expect the energy to not depend on the radius $R$ and, more importantly, to be linear with the angle deficit $\Delta{\phi}$.

### v2

# On Nanocones as Gravitational Analog Systems

@@ Abstract
This study investigates the fundamental properties of graphene and boron nitride (BN) nanostructures, exploring their torsional energy characteristics within the framework of Teleparallel Equivalent of General Relativity (TEGR). By constructing nanocones with disclination defects in these materials, the linear dependence of torsional energy on the disclination angle is analyzed, as predicted by TEGR. The qualitative validation of TEGR’s energy expression is supported by simulations, which show a strong correlation between torsional energy and the disclination angle, consistent with theoretical predictions. Additionally, a quantitative analysis is proposed by estimating the coupling constant $\kappa$ associated with TEGR through molecular simulations and Density Functional Theory (DFT) calculations. The results suggest that $\kappa$ reflects the interatomic forces within the materials, providing insights into the nature of spacetime and gravitational interactions on a microscopic scale. These findings contribute to the understanding of material physics and offer implications for the precision and validity of TEGR in describing gravitational phenomena.
Keywords: Teleparallel Equivalent of General Relativity, topological defects, geometric theory of defects, Universal Force Field, nanocones.
March 27, 2025
@@ I Introduction
Analog models in physics have long existed. A mathematical equivalence between the equations governing certain physical phenomena allows for analogy between them. A good example of this occurred in the realm of electrical circuits. In the early 19th century, when instrumentation for measuring electrical and magnetic quantities was being developed, it was very costly to conduct investigations in circuits containing capacitors and inductors. Therefore, mechanical analog models were employed, which facilitated a better understanding of electrical circuits, as well as simpler experimentation. Analog models thus enable the testing of physical concepts without the need to access the physical system of interest. Experimental black hole physics is still in its infancy and could greatly benefit from the use of gravitational analog models barcelo . While experimenting with gravitational waves can shed light on certain aspects of black hole physics, testing other concepts such as Hawking radiation requires the use of analog models, often established in condensed matter. However, there are also acoustic analogs; acoustic black holes, proposed by Unruh Unruh , have been used to understand how event horizons can produce a spectrum of thermal radiation Unruh1 . Hence, gravitational analog models have become important tools for understanding black hole physics. One such example was the construction of an experimental apparatus where a robophysical model was built to simulate the motion of particles around a Schwarzschild singularity li2023robophysical , an experiment that culminated in very satisfactory results compared to the expected values from GR (General Gelativity).
@@
GR is the most widely accepted theory in the literature describing gravitation. However, it is fraught with inconsistencies. Perhaps the most important of these is the definition of gravitational energy. Indeed, there are several attempts to define such an important concept in physics within the scope of general relativity, but none of them meets the necessary requirements for a consistent definition, namely: i) the energy should be independent of the coordinate system used; ii) it should be the component of a covariant four-vector under global Lorentz transformations and therefore should depend on the choice of reference frame. Alternatively, there are approaches that describe the gravitational field, in which the concept of energy is more natural. In this article, an alternative theory dynamically equivalent to general relativity, called Teleparallel Equivalent to General Relativity (TEGR) maluf , will be used. In the context of TEGR, there exists a consistent energy-momentum four-vector, so this quantity can be used to experimentally determine which approach correctly describes gravitation. In other words, from a dynamic point of view, both GR and TEGR predict the same behaviors. On the other hand, a consistent definition of energy exists only in TEGR (for a complete discussion of gravitational energy, we encourage the reader to consult Ref. annalen ). On one hand, there are currently no experiments designed to deal with gravitational energy. On the other hand, the use of gravitational analogue models may lead to clues suggesting the supremacy of TEGR over GR, if gravitational energy is confirmed in such systems.
@@
Graphene is a material predicted in the mid-20th century and synthesized in the early 21st century grafeno . It consists of a two-dimensional lattice of carbon atoms arranged in a hexagonal shape, giving graphene highly desired properties such as good conductivity and strength. The equation governing the dynamics of electrons in graphene is relativistic, as the Dirac equation embodies Lorentz transformations. The Dirac equation also naturally couples to the dynamic variables of TEGR, making graphene an excellent analog gravitational system that could allow investigation into the physical reality of gravitational energy.
Recent studies have explored alternative approaches to describing the electronic and geometric properties of graphene. In particular, metric-based methods have been employed to investigate graphene wormholes and their potential for superconductivity. For instance, the work of Capozziello et al. formulates an effective geometric framework in which superconducting properties emerge from the topology of graphene sheets connected via Chern-Simons bridges referee1 . Similarly, Ref. referee2 examines the role of modified gravity-inspired geometric structures in electronic transport within graphene, drawing analogies between crystalline defects and spacetime modifications. These approaches have shown the relevance of curvature-based descriptions in modeling the electronic properties of graphene.
@@
In pursuit of our goal, we consider a graphene sheet as a background flat geometry and impose a gravitational field by inserting a topological defect into the sheet. We choose a disclination defect to construct a well-known nanostructure for graphene, i.e., a nanocone charlier2001electronic ; adhikari2012stabilities ; ardeshana2017approach . Consequently, we evaluate the gravitational energy of the disclination using TEGR and compare the predicted qualitative behavior of the torsional energy with the energy associated with the torsion potential obtained from computational simulations within Universal Force Field (UFF) . The UFF is a versatile molecular mechanics force field used in computational chemistry. It employs empirical parameters to model interatomic interactions, including bonded and non-bonded terms, torsional potentials, and geometric optimizations. UFF is widely applied in molecular simulations for accurate predictions of molecular structures, energies, and properties. The main aspects of this theory necessary for the comprehension of the simulations will be presented in this article. We also infer about the possibility of the physical origin of the coupling constant in graphene and obtain a numerical value for the energy within the order of magnitude given by the simulations. Therefore, we aim to address these aspects in this article to support future experiments designed to measure gravitational energy, especially since the synthesis of nanocones is already a reality experimental .
@@ II Teleparallelism Equivalent to General Relativity (TEGR)
TEGR is a geometrical theory of gravity that was first considered by Einstein in an attempt to unify gravity and electromagnetism. Several decades after being abandoned by Einstein, Møller revived the idea with another objective: to establish a true energy-momentum for the gravitational field moller1964conservation . Despite the many decades of GR, researchers have not been able to define a true energy-momentum tensor for the gravitational field, with several pseudo-tensors defined in the literature, all of which depend on the choice of coordinates. Therefore, a definition for the energy of a gravitational field is only possible globally, i.e., by integrating the pseudo-tensors across the entire spacetime. Such integration is possible only if the spacetime is asymptotically flat at spatial infinity. After a few years of attempting to define a true energy-momentum tensor for the gravitational field, Møller realized that the main problem with GR was the fact that its fundamental variable is the metric tensor $g_{\mu\nu}$ and concluded that he could solve the problem by formulating GR with the tetrad field $e_{a\mu}$ as its fundamental variable. Furthermore, Maluf realized that the field equations of the theory allow for a definition of a true energy-momentum tensor for the gravitational field. In the following, we follow the development of TEGR as presented in Ref. maluf .
@@
The geometry in which TEGR is established is the Weitzenböck geometry. Given the prominence in the literature of topological defects considered within Riemann-Cartan geometry, we depart from Riemann-Cartan to the Weitzenböck one, where the Levi-Civita connection is responsible for the rotation symmetry instead of the affine connection in Riemann-Cartan geometry. Therefore, we shall show that TEGR admits linear rotations (disclinations) and translations (dislocations) as topological defects in its geometry, similar to Riemann-Cartan.
In a manifold endowed with metricity, the structure equations can be obtained from the co-frame 1-form $e^{a}=e^{a}\,_{\mu}dx^{\mu}$ and the connection 1-form $\omega^{a}\,_{b}=\omega^{a}\,_{b\mu}dx^{\mu}$, where $\omega^{a}\,_{b\mu}$ are the 0-form components of the connection. The torsion 2-form $T^{a}$ of the manifold can be obtained by the exterior total derivative $D$ of the co-vector, i.e.,
@@
$=\frac{1}{2}T^{a}\,_{\mu\nu}dx^{\mu}\wedge dx^{\nu}$ (2)
@@
$T^{a}\,_{\mu\nu}=\partial_{\mu}e^{a}\,_{\nu}-\partial_{\nu}e^{a}\,_{\mu}+\omega^{a}\,_{b\mu}e^{b}\,_{\nu}-\omega^{a}\,_{b\nu}e^{b}\,_{\mu}$ (3)
@@
$=\frac{1}{2}R^{a}\,_{b\mu\nu}dx^{\mu}\wedge dx^{\nu}\,,$ (4)
@@
$R^{a}\,_{b\mu\nu}$ $=\partial_{\mu}\omega^{a}\,_{b\nu}-\partial_{\nu}\omega^{a}\,_{b\mu}$
$+\omega^{a}\,_{c\mu}\omega^{c}\,_{b\nu}-\omega^{a}\,_{c\nu}\omega^{c}\,_{b\mu}$ (5)
@@
$\nabla_{\mu}e_{a\nu}=\partial_{\mu}e_{a\nu}-\Gamma^{\lambda}\,_{\mu\nu}e_{a\lambda}=0\,.$ (6)
@@
$\Gamma^{\lambda}\,_{\mu\nu}=e^{a\lambda}\partial_{\mu}e_{a\nu}$ (7)
@@
$T^{a}\,_{\mu\nu}=\partial_{\mu}e^{a}\,_{\nu}-\partial_{\nu}e^{a}\,_{\mu}\,.$ (8)
@@
$K_{\lambda\mu\nu}$ $=$ $\frac{1}{2}(T_{\lambda\mu\nu}+T_{\mu\lambda\nu}-T_{\nu\lambda\mu})\,,$ (9)
can be evaluated. Since the affine connection is chosen to be zero, the components of the Levi-Civita connection $\mathring{\omega}_{ab\mu}$ can be evaluated using the relation
$\omega_{ab\mu}=\mathring{\omega}_{ab\mu}+K_{ab\mu}=0$ (10)
@@
$\mathfrak{L}(e_{a\mu})$ $=$ $-\kappa\,e\Sigma^{abc}T_{abc}-\frac{1}{c}\mathfrak{L}_{M}$ (12)
@@
-- $\frac{1}{c}\mathfrak{L}_{M}\,,$
where $e=det(e^{a}\,_{\mu})$, $\kappa=c^{3}/16\pi G$ in SI units, and $\mathfrak{L}_{M}$ stands for the matter-radiation fields Lagrangian density. The torsion invariant, and hence the Lagrangian density, is equal to the curvature scalar constructed from the Levi-Civita connection plus a total divergence term, i.e.,
@@
Although the Lagrangian density (12) is not invariant under local Lorentz transformations, the field equations (14) are covariant under coordinate transformations and under the full group of Lorentz transformations. It can be shown that (14) may be written as $R_{a\mu}-\frac{1}{2}e_{a\mu}R=\frac{1}{2c\kappa}T_{a\mu}$, i.e., they are equivalent to Einstein’s equations. Therefore, TEGR yields the same dynamics as GR but allows for the definition of a true energy-momentum tensor.
@@ III Topological Defects in spacetime
The matrix $\mathcal{G}^{(i)(j)}$, called Frank matrix, represents the rotation matrix along the axis containing the angle deficit, i.e., the rotation angle represents the amount that we must rotate to perform a full $2\pi$ lap in spacetime. The curvature tensor of the affine connection is zero for the Weitzenböck geometry. However, the curvature tensor of the Levi-Civita connection $\mathring{\omega}_{ab\mu}$ is nonzero in the general case. This curvature tensor allows for a consistent definition of disclination in TEGR. This feature will be demonstrated here for the defect considered.
A conical defect along the $z$ direction is characterized by the disclination angle parameter $\Delta\phi=2\pi\big{(}1-\beta(r)\big{)}$ in the line element
@@
When $\beta=1$, we recover Minkowski spacetime. The disclination angle can be evaluated from (24) by computing the length of the circumference along the $z$ axis, but we will consider relations (22) and (23). First, we need a set of tetrads associated with the metric (24) and adapted to a static observer. One set satisfying the condition is given by
@@
$\mathring{\omega}^{(1)(2)}\,_{2}=1-\partial_{r}(r\,\beta)\,.$ (27)
@@
i.e., the rotation matrix of the disclination angle $\Delta\phi=2\pi(1-\beta_{0})$ along the $z$ axis. Thus, TEGR allows torsion and curvature and is consistent in computing topological defects in spacetime, with the rotational defects given by the Levi-Civita connection constructed out of the contortion tensor given by (10).
@@
Hence, the zeroth component of $P^{a}$ contained within a cylinder of radius $R>a$ and length $L$ is given by
$P^{(0)}$ $=2\kappa\int_{V}\partial_{r}\big{(}1-\partial_{r}(r\beta)\big{)}dV$
@@
$=4\pi\,\kappa\,L(1-\beta_{0})$
$=\frac{c^{3}}{8\pi G}\,L\,\Delta\phi\,.$ (39)
From (39), we see that with TEGR, we expect the energy $E=cP^{(0)}$ to not depend on the radius $R$ and, more importantly, to be linear with the angle deficit $\Delta{\phi}$.